OpenStudy (anonymous):

I need help please! Suppose a computer that costs $3000 new is only worth $600 after 3 years .What is the average annual rate of depreciation?

OpenStudy (amistre64):

600 = 3000(1 - r)^3

OpenStudy (amistre64):

600/3000 = (1-r)^3 cbrt(6/30) = 1-r r = 1 - cbrt(5)

OpenStudy (amistre64):

...... r = 1 - cbrt(1/5) might be better ;)

OpenStudy (anonymous):

what is answer?

OpenStudy (anonymous):

cbrt?

OpenStudy (amistre64):

cbrt(....) means cube root

OpenStudy (anonymous):

than What is the average annual rate of depreciation?

OpenStudy (amistre64):

it depreciates by 41.5 % each year

OpenStudy (anonymous):

\[1-\sqrt[3]{\frac{1}{5}} ?\]

OpenStudy (amistre64):

well that dint format right

OpenStudy (amistre64):

\[rate=1-\sqrt[3]{1\over5}=abt. \ 41.5\%\]

OpenStudy (anonymous):

Ok, thank

OpenStudy (anonymous):

Considering composed taxes: \[600=3000*(1-i)^3\] \[\log(6/30)=3*\log(1-i)\] \[10^{\frac{\log(6/30)}{3}}={1-i}\] \[1-10^\frac{\log(6/30)}{3}=i\] \[i \approx 0,4152 \approx 41,52\% \]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

thank LuisGuihermepegas I like the way you write